help plzzProblem 1-19 Traditional and Contribution Format Income Statements (LO1-6] Todrick Company is a merchandiser that reported th
Answer #1

Solution

Before giving solutions to other parts, value of the missing parts are to be determined,

At first, we know that,

(Contribution - Total Fixed Costs) = Net Operating Income

So, Let us replace the values as per problem,

(Contribution - FIxed Administrative Expenses - Fixed Selling Expenses) = Net Operating Income

Or, 90000 - 18000 - Fixed Selling Expenses = 27000

Or, Fixed Selling Expenses = 90000 - 18000 - 27000

Or, Fixed Selling Expenses = 45000

Therefore, Fixed Selling Expenses = $ 45,000

Again,

(Sales - Variable Costs) = Contribution

Or, Sales - Manufacturing Expenses - Variable Selling Expenses - Variable Administrative Expenses = Contribution

Or, Sales - (Opening Inventory + Purchases - Closing Inventory) - Variable Selling Expenses - Variable Administrative Expenses = Contribution

Or, 450000 - (30000 + 300000 - 15000) - 22000 - Variable Administrative Expenses = 90000

Or, 450000 - 315000 - 22000 - Variable Administrative Expenses = 90000

Or, Variable Administrative Expenses = 450000 - 315000 - 22000 - 90000

Or, Variable Administrative Expenses = 22500

Therefore, Variable Administrative Expenses = $ 22,500

(1)

Todrick Company
Contribution Format Income Statement
Particulars Amount ($) Amount ($)
Sales 4,50,000.00
Less: Variable Expenses
1. Manufacturing Expenses
Purchases of Merchandise 3,00,000.00
Add: Beginning Merchandise Inventory      30,000.00
Less: Ending Merchandise Inventory    -15,000.00 3,15,000.00
2. Variable Selling Expenses      22,500.00
3. Variable Administrative Expenses      22,500.00 3,60,000.00
Contribution      90,000.00
Less: Fixed Expenses
1. Fixed Selling Expenses      45,000.00
2. Fixed Administrative Expenses      18,000.00      63,000.00
Net Operating Income      27,000.00

(2)

Todrick Company
Traditional Format Income Statement
Particulars Amount ($) Amount ($)
Sales    4,50,000.00
Less: Cost of Goods Sold
Beginning Merchandise Inventory      30,000.00
Add: Purchases of Merchandise 3,00,000.00
Less: Ending Merchandise Inventory    -15,000.00 -3,15,000.00
Gross Profit    1,35,000.00
Less: Operating Expenses
Variable Selling Expenses      22,500.00
Variable Administrative Expenses      22,500.00
Fixed Selling Expenses      45,000.00
Fixed Administrative Expenses      18,000.00 -1,08,000.00
Net Operating Income        27,000.00

(3)

Total Sales = $ 450,000

Number of Units = 1,000

Selling Price per unit = $ (450,000 / 1,000) = $ 450

Therefore, Selling Price per unit = $ 450 per unit

(4)

Total Variable Cost (Refer Contribution Format Income Statement) = $ 360,000

Number of Units = 1,000

Variable Cost per unit = $ (360,000 / 1,000) = $ 360

Therefore, Variable Cost per unit = $ 360 per unit

(5)

Contribution Margin = $ 90,000

Number of Units = 1,000

Contribution per unit = $ (90,000 / 1,000) = $ 90

Therefore, Contribution per unit = $ 90 per unit

(6)

It is suggested to go for Contribution Format Income Statement. Although Traditional Format can also give an overview about the sensitivity of Net Operating Income with Unit Sales, but it is to be remembered that Net Operating Income is also influenced by Variable Costs. If there is a change in unit sales, it shall also affect the Variable Costs. Hence it is neither sufficient nor practicable to measure Profit sensitivity on the basis of the sales as variable costs should also be considered for the measurement purpose.